# Normalized Matching Transformer

### v1

@@ Abstract
We present a new state of the art approach for sparse keypoint matching between pairs of images. Our method consists of a fully deep learning based approach combining a visual backbone coupled with a SplineCNN graph neural network for feature processing and a normalized transformer decoder for decoding keypoint correspondences together with the Sinkhorn algorithm. Our method is trained using a contrastive and a hyperspherical loss for better feature representations. We additionally use data augmentation during training.
This comparatively simple architecture combining extensive normalization and advanced losses outperforms current state of the art approaches on PascalVOC and SPair-71k datasets by $5.1\%$ and $2.2\%$ respectively compared to BBGM[31], ASAR [30], COMMON [20] and GMTR [33] while training for at least $1.7x$ fewer epochs.
@@ 1 Introduction
While pure deep learning methods have already reached very high results on keypoint matching datasets, we show that there is still room for improvement. First, better backbones boost performance. We replace the commonly used VGG [33] backbone for a current swin-transformer [22]. Second, we process features at keypoints with a SplineCNN GNN [10], which adds helpful inductive biases incorporating the geometry of the keypoints to match. Third, we use transformers to mix information between and across images. We show that vanilla transformers can be outperformed by additional normalization techniques used in normalized transformers [23]. We argue that the employed normalization techniques are well-suited for our normalized feature representation: Instead of only normalizing at the end before doing cosine similarity and computing the losses, we normalize throughout, which helps in faster training and better overall performance. Our pipeline is trained using a contrastive [27] and hyperspherical [25] loss together with data augmentation.
@@ 3 Method
In particular, the normalized transformer uses normalization after attention, the MLP block and the residual connections. Let $f=f_{1},\ldots,f_{m}$ be a feature sequence coming from all the keypoints in an image. For cross-attention let $f_{other}$ be the keypoint feature sequence from the other image. Then normalized self-attention, cross-attention and MLP layers can be written as
@@
$f_{A}\leftarrow\text{Norm}(\text{Cross-Attn}(f,f_{other}))$;
@@
 2 Norm. Cross-Attn($f,f_{other}$):
@@ 4 Normalized Matching Transformer
$\mathcal{L}_{\mathbf{HS}}=\frac{1}{m}\sum_{i=1}^{m}\text{max}_{j}\left\{C_{ij}-2\cdot\mathbf{1}_{\{i=j\}}\right\}\,.$ (3)
This loss penalizes whenever two different keypoints are aligned. The diagonal subtraction $-2\cdot\mathbf{1}_{\{i=j\}}$ ensures that this penalization is not carried out for the same keypoint.
@@ 4 Experiments
Our network is trained using the Adam optimizer [18]. The initial learning rate for the network is set to $5\times 10^{-4}$, while the swin-transformer [22] backbone with layer normalization uses a learning rate scaled down by a factor of $0.03$. Additional normalization was omitted due to the inherent design of the normalized transformer. The learning rate is scaled by a factor of $0.1$ after epoch 2 and 5. For PascalVOC the batch size of image pairs is set to 8 and for Spair-71k the batch size is set to 5. Our validation set is obtained by taking $1000$ image pairs per class. We use augmentations provided through the Albumentations package [5]. Specifically, we use Mixup [45], Cutmix [43] and Random Erasing [47]. The model is run on two GPUs by using torch’s Distributed library. Other hyperparameters are listed in Table 1.
@@
Our normalized matching transformer approach as detailed in Algorithm 4.
@@ Results.
We provide ablations on our main architectural and training contributions on PascalVOC. The results are summarised in Table 4. We see that all contributions significantly add to final performance. The main performance driver are our losses, followed by improved backbone, normalized transformer and augmentations. Augmentations, while giving $1.2\%$, are still significant, but not overly so. We have experimented with other simple pixel-wise augmentations like changing saturation value, random gamma, RGB shift etc., which however degraded performance.
@@ 5 Conclusion
We have shown that the combination of advances in transformer design coupled with representation learning losses leads to significant gains in two very competitive keypoint matching benchmarks. We argue that normalization throughout our neural networks is helpful in the context of keypoint matching, where typically only the output features have been normalized. This has lead to fast training and superior performance. We conjecture that pursuing normalization together with representation learning in other related computer vision applications might also lead to improvements.

### v2

@@ Abstract
We present a new state of the art approach for sparse keypoint matching between pairs of images. Our method consists of a fully deep learning based approach combining a visual backbone coupled with a SplineCNN graph neural network for feature processing and a normalized transformer decoder for decoding keypoint correspondences together with the Sinkhorn algorithm. Our method is trained using a contrastive and a hyperspherical loss for better feature representations. We additionally use data augmentation during training. This comparatively simple architecture combining extensive normalization and advanced losses outperforms current state of the art approaches on PascalVOC and SPair-71k datasets by $5.1\%$ and $2.2\%$ respectively compared to BBGM [31], ASAR [30], COMMON [20] and GMTR [13] while training for at least $1.7x$ fewer epochs.
@@ 1 Introduction
While pure deep learning methods have already reached very high results on keypoint matching datasets, we show that there is still room for improvement. First, better backbones boost performance. We replace the commonly used VGG [33] backbone for a current swin-transformer [22]. Second, we process features at keypoints with a SplineCNN GNN [10], which adds helpful inductive biases incorporating the geometry of the keypoints to match. Third, we use transformers to mix information between and across images. We show that vanilla transformers can be outperformed by additional normalization techniques used in normalized transformers [23]. We argue that the employed normalization techniques are well-suited for our normalized feature representation: Instead of only normalizing at the end before doing cosine similarity and computing the losses, we normalize throughout in the normalized transformer, which helps in faster training and better overall performance. Our pipeline is trained using a contrastive [27] and hyperspherical [25] loss together with data augmentation.
@@ 3 Method
In particular, the normalized transformer uses normalization after attention, the MLP block and the residual connections. Let $f=f_{1},\ldots,f_{m}$ be a feature sequence coming from all the keypoints in an image. For cross-attention let $f_{other}$ be the keypoint feature sequence from the other image. Then normalized self-attention, cross-attention and MLP layers can be written as:
@@
$f_{A}\leftarrow\text{Norm}(\text{Cross-Attn}(f,f_{o}))$;
@@
 2 Norm. Cross-Attn($f,f_{o}$):
@@ 4 Normalized Matching Transformer
$\mathcal{L}_{\mathbf{HS}}=\sum_{j=1}^{n}\max_{j\neq i}C_{ij}\,.$ (3)
This loss penalizes whenever two different keypoints are aligned. The $\max_{j\neq i}$ ensures that this penalization is not carried out for the same keypoint.
@@ 4 Experiments
Our network is trained using the Adam optimizer [18]. The initial learning rate for the network is set to $5\times 10^{-4}$, while the swin-transformer [22] backbone with layer normalization uses a learning rate scaled down by a factor of $0.03$. Additional normalization was omitted due to the inherent design of the normalized transformer. The learning rate is scaled by a factor of $0.1$ after epoch 2 and 5. For PascalVOC the batch size of image pairs is set to 8 and for Spair-71k the batch size is set to 5. Our validation set is obtained by taking $1000$ image pairs per class. We use augmentations provided through the Albumentations package [5]. Specifically, we use Mixup [45], Cutmix [43] and Random Erasing [47]. The model is trained on one A100 GPU with 40 GB memory and takes about 9 hours for PascalVOC and 7 hours for SPair-71k. Other hyperparameters are listed in Table 1.
@@
Our normalized matching transformer approach as detailed in Algorithm.
@@ Results.
Inference Speed
We measure inference speed on a single NVIDIA GeForce RTX 4090 (batch size = 1), averaging over 1000 forward passes. The complete Normalized Matching Transformer processes each image pair in 44.4 ms, with the backbone + SplineCNN accounting for 39 ms and the two decoders for 5.4 ms.
@@
We provide ablations on our main architectural and training contributions on PascalVOC. The results are summarised in Table 4. We see that all contributions significantly add to final performance. The main performance driver are our losses, followed by improved backbone, normalized transformer and augmentations. Augmentations, while giving $1.2\%$, are still significant, but not overly so. The backbone ablation shows that we obtain better results on PascalVOC even when we employ the VGG backbone (i.e. 83.8%) also used by COMMON [20], ASAR [30] and BBGM [31] and we even slightly outperform GMTR [13] which uses a swin-transformer backbone.
We have also experimented with other simple pixel-wise augmentations like changing saturation value, random gamma, RGB shift etc., which however degraded performance.
@@ 5 Conclusion
We have shown that the combination of advances in transformer design coupled with representation learning losses leads to significant gains in two very competitive keypoint matching benchmarks. We argue that using normalization throughout our neural networks and in the losses is helpful in the context of keypoint matching, where in previous work at most the output features have been normalized. This joint normalization leads to fast training and superior performance. We conjecture that pursuing normalization together with representation learning in other related computer vision applications might also lead to improvements.
6 Broader Impact
Our work improves accuracy of keypoint matching. Since other methods for this task already exist that work adequately, we do not foresee any societal ramifications that go beyond existing systems.